how many cubic meters are there in 7000cm cube
Answer by mananth(16949)

how many cubic meters are there in 7000cm cube
1meter =100 cms
1 cubic meter = 100*100*100 cm^3
=1,000,000
7000 / 1000000= 7/1000 meter cube
0.007 cu. meter